# The differences of baryonic and dark matter scaling relations in galaxy clusters between IllustrisTNG simulations and observations

### v1

@@
The differences of baryonic and dark matter scaling relations in galaxy clusters between IllustrisTNG simulations and observations
@@ 1 Introduction
One of the last century’s key discoveries was that the Universe is expanding. Nowadays, we have a very well-constrained “standard model”, called ”Lambda Cold Dark Matter” ($\Lambda$CDM), which says that the Universe is composed of dark energy ($\Omega_{\Lambda}\sim 0.69$) and matter ($\Omega_{\rm m}=\Omega_{\rm b}+\Omega_{\rm dm}\sim 0.31$), which $\Omega_{\rm b}$ represents the baryonic density and $\Omega_{\rm dm}$ the dark matter density. From the matter density, $15.7$ per cent corresponds to baryons, while $\sim$ 84.3 per cent is attributed to dark matter (Planck Collaboration et al., 2016a). This predominance of dark matter over visible matter significantly impacts the gravitational processes driving the formation of structures; briefly, after the Big Bang, small quantum perturbations in the density field lead to matter overdensities. As Dark Matter only interacts gravitationally, it had time to clump up and grow, forming the first dark matter haloes. These regions started to grow by accreting baryons and smaller haloes (White & Rees, 1978; White & Frenk, 1991). Several authors have thoroughly reviewed this ”hierarchical structure formation” scenario in the last decades through analytical studies (e.g. Press & Schechter, 1974; White & Rees, 1978; White & Frenk, 1991) and numerical simulations (e.g. White, 1978; Villumsen, 1982; McGee et al., 2009; Pallero et al., 2019).
@@
Within these haloes, galaxy clusters are a key piece in the puzzle to understand the evolution of the Universe, as they are the largest gravitationally bound structures. To form this kind of structure, matter of large regions of the Universe should collapse ($\sim 10$Mpc); so they are expected to have a baryon fraction close to the universal value, $f_{\rm bar}^{\rm U}\sim\Omega_{\rm b}/\Omega_{\rm m}\sim 0.157\pm 0.004$ (Planck Collaboration et al., 2016a). Observational results indicate that massive clusters $\log_{10}{\left(M_{\rm 200c}/\rm M_{\odot}\right)}\gtrsim 14.5$ typically exhibit a baryonic fraction close to this universal value, as expected (Giodini et al., 2009). Furthermore, Sunyaev-Zel’dovich (SZ) surveys and X-ray measurements have confirmed similar results (Chiu et al., 2018, C18 hereafter). However, low-mass clusters and groups of galaxies do not fulfill these expectations (e.g. Ettori, 2003; McGaugh et al., 2010; Lovisari, L. et al., 2015; Chiu et al., 2018).
@@
The last point highlights a discrepancy between early-time observations of baryon fractions in the Cosmic Microwave Background (CMB) (e.g. Planck Collaboration et al., 2016a) and late-time Sunyaev-Zel’dovich (SZ) and X-ray observations of galaxy groups and clusters. For example, C18 analyzed the baryon fraction, $f_{\rm bar,500c}=f_{\rm star,500c}+f_{\rm hot-gas,500c}$, of 91 self-similar clusters in the redshift range $1.250.20<z<$ 1.25, using data from the SPT-SZ survey and Chandra X-ray observations. By employing scaling relations for X-ray hot gas and galaxy stellar mass, they found that clusters with $\log_{10}{\left(M_{\rm 500c}/\rm M_{\odot}\right)}\gtrsim 14.9$ reach a baryon fraction close to the universal value. However, clusters with $\log_{10}{\left(M_{\rm 500c}/\rm M_{\odot}\right)}\lesssim 14.9$ exhibit significantly lower baryon fractions. These observations suggest that not all baryonic matter in low-mass clusters is detected, even though it should reside within their halo boundaries (Shull et al., 2012; Sanderson et al., 2013). This discrepancy presents a key challenge in the $\Lambda$CDM model, commonly called the ”missing baryon” problem. Additionally, C18 found that baryon fractions vary strongly with total mass but weakly with redshift. These two properties are inconsistent with a hierarchical scenario, such as the most massive clusters, with baryon fractions near the universal value, could not have formed through the accretion of low-mass clusters with lower baryon fractions.
@@
Studies from cosmological simulations and observations led to two main explanations for the missing baryons. The first suggests the existence of warm-hot gas in the intergalactic medium (WHIM) at temperatures of $10^{5}-10^{7}$ K (Cen & Ostriker, 1999). Nicastro et al. (2018) detected diffuse WHIM using two OVII absorption line tracers, although their measurements involve considerable uncertainties, as the signals could originate from sources other than WHIM itself (Johnson et al., 2019). Additionally, de Graaff et al. (2019) detected WHIM through the Planck thermal Sunyaev-Zel’dovich (tSZ) signal (Planck Collaboration et al., 2016b), and WHIM has been characterized and analyzed in magnetohydrodynamical simulations (e.g. Martizzi et al., 2019). The second explanation posits that the missing baryons reside in regions outside the halo boundary, such as $R_{500\rm c}$ or $R_{200\rm c}$ (e.g. Haider et al., 2016; Ayromlou et al., 2021; Gouin et al., 2022; Ayromlou et al., 2023). Feedback processes redistribute baryons to the outskirts of the halo, particularly in haloes with $\log_{10}\left(M_{\rm 200c}/M_{\odot}\right)\lesssim 14$ (Ayromlou et al., 2023, A23 hereafter). Using simulations, A23 introduced the closure radius, $R_{\rm c}$, as the radius where the baryon fraction approaches the Universal value. For example, for haloes with $\log_{10}\left(M_{\rm 200c}/M_{\odot}\right)\lesssim 14$ , AGN feedback processes redistribute baryons up to $R_{\rm c}\sim 1.5-2.5R_{\rm 200c}$ (Ayromlou et al., 2023). These two explanations are not mutually exclusive and provide complementary insights into galaxy evolution and large-scale structure cosmology (Walker et al., 2019; Nicastro et al., 2022).
@@
The scaling relations are powerful tools for investigating the power laws associated with observables in haloes. The self-similar model (Kaiser, 1986) provides a mathematical framework for describing scaling relations in hydrostatic equilibrium structures. These tools allow the connection of large samples of galaxy clusters to the measurement of cosmological parameters (Haiman et al., 2001; Holder et al., 2001; Carlstrom et al., 2002). Scaling relations have been extensively studied through X-ray observations (e.g. Mohr & Evrard, 1997; Arnaud & Evrard, 1999; Mohr et al., 1999; Reiprich & Böhringer, 2002; O’Hara et al., 2006; Arnaud et al., 2007; Pratt et al., 2009; Sun et al., 2009; Vikhlinin et al., 2009; Mantz et al., 2016; Chiu et al., 2018) and hydrodynamical simulations (e.g. Evrard, 1997; Bryan & Norman, 1998; Nagai et al., 2007; Stanek et al., 2010; Pillepich et al., 2018; Pop et al., 2022a; Hadzhiyska et al., 2023a). The inclusion of AGN feedback processes in simulations significantly affects the X-ray and SZ scaling relations in galaxy groups (Puchwein et al., 2008; Fabjan et al., 2011; Pike et al., 2014; Planelles et al., 2013; Le Brun et al., 2016; Truong et al., 2017; Henden et al., 2018, 2019; Lim et al., 2021; Yang et al., 2022). Pop et al. (2022a, P22 hereafter) introduced a novel method for fitting a scaling relation using a General Double Power Law (GDPL), model derived from the broken power law (BPL: Jóhannesson et al., 2006) scheme. The GDPL model demonstrated superior performance in representing self-similarity for both relaxed and unrelaxed haloes at $\log_{10}\left(M_{\rm 500c}/{\rm M_{\odot}}\right)\gtrsim 14$ compared to the simple power law (SPL: Lotka, 1926) in simulations. If the trends and slopes of the SPL scaling relations presented by C18 can be reproduced in simulations, the missing baryons can be identified and further investigated.
@@
In this work, we compare the self-similar scaling relations of galaxy clusters obtained from the observations of C18 with the hydrodynamical simulations. We use the suite of the cosmological magnetohydrodynamical simulations from the IllustrisTNG project, selecting a sample of haloes with $M_{\rm 200c}\geq 7\times 10^{13}$ $\rm M_{\odot}$ to match the observational mass conditions and scaling relations of C18. To study their behavior separately, cold and warm gas components are introduced into these scaling relations. We analyze the percentage distribution of the baryonic components of the haloes: cold gas, WHIM, hot gas, stellar mass from galaxies, and intracluster light (ICL). Finally, we fit the SPL, BPL, and GDPL models and examine their respective shapes.
@@
This paper is structured as follows. In Section 2, we describe the methods used in this research, including simulations and observational data. Section 3 outlines the formalism of the baryonic self-similar scaling relations, incorporating both the SPL and GDPL approaches. In Section 4, we present the results of the SPL, BPL, and GDPL models for the baryonic fraction scaling relations, comparing these models with observational data, varying the galaxy aperture, and adding the remaining gas components. In Section 5, we discuss the implications of these results, particularly in the context of baryonic component accretion. Finally, in Section 6, we summarise our findings.
@@ 2 Simulations and observational data
2.1 IllustrisTNG: the state-of-art magneto-hydrodynamical and cosmological simulations
IllustrisTNG (Naiman et al., 2018, Nelson et al., 2018, Springel et al., 2018, Pillepich et al., 2018, Marinacci et al., 2018) is a comprehensive suite of state-of-the-art magnetohydrodynamical and cosmological simulations focused on the formation and evolution of galaxies. It uses the moving-mesh AREPO code (Springel, 2010) to solve the Poisson’s equations. Additionally, it solves magnetohydrodynamic equations using a Voronoi tessellation method (Pakmor et al., 2011; Pakmor & Springel, 2013). It includes a physical model for galaxy formation, incorporating non-gravitational processes such as gas radiative cooling, star formation, stellar evolution, and stellar feedback (Pillepich et al., 2018). Additionally, IllustrisTNG accounts for the physics of supermassive black holes, including their seeding, merging, and feedback processes (Weinberger et al., 2017).
@@ 2.1 IllustrisTNG: the state-of-art magneto-hydrodynamical and cosmological simulations
We define galaxy clusters as self-gravitating haloes with masses $M_{\rm 200c}\geq 7\times 10^{13}$ $\rm M_{\odot}$ (for any redshift), comparable to the mass of the Fornax cluster (Drinkwater et al., 2001). The haloes and subhaloes were identified in a two-step fashion. First, haloes were selected using the Friends-of-Friends algorithm (FoF: Davis & Djorgovski, 1985; Springel et al., 2001), with a linking length of $b=0.2$, applied only to dark matter particles. Then, subhaloes were selected within FoF haloes using the subfind algorithm (Springel et al., 2001; Dolag et al., 2009), as those self-gravitating structures residing within haloes. In this work, we define galaxies as those subhaloes which contain 1000 stellar particles, i.e., $M_{\rm star}\geq 5\times 10^{9}$ $\rm M_{\odot}$ for TNG100 and $M_{\rm star}\geq 5\times 10^{10}$ $\rm M_{\odot}$ for TNG300, to ensure enough resolution to avoid spurious identifications. Gas within structures is split into three components: cold gas ($T<10^{5}$ $\rm K$), warm gas ($10^{5}\leq T<10^{7}$ $\rm K$), and hot gas ($T\geq 10^{7}$ $\rm K$) (Cen & Ostriker, 1999; Martizzi et al., 2019; Gouin et al., 2022, 2023). The selection of these temperature thresholds stems from the fact that the star formation model in IllustrisTNG is based on an effective equation of state and a representation of cold and warm gas phases within each cell (Pillepich et al., 2018). Moreover, these thresholds are widely employed in IllustrisTNG studies to distinguish gas phases based on temperature and electron density (e.g. Martizzi et al., 2019; Gouin et al., 2022, 2023), as they effectively capture the gas phase diagram. Clusters were selected within $0.0\leq z\leq 1.5$, with a time window between snapshots of 0.5 Gyr (TNG100) and 1.8 Gyr (TNG300). This resulted in a sample of 218 and 1605 clusters respectively.
@@
For each cluster, we define a radius $R_{\Delta}=R_{\rm 500c}$, where the density is $500$ times the critical density of the Universe (see Section 3). Then the baryon (gas and stars), the stellar, and the hot, warm and cold gas masses are all defined by adding the corresponding particles with $R_{\Delta}$. The total mass, $M_{\rm 500c}$, is the sum of $M_{\rm bar,500c}$ and the dark matter mass, $M_{\rm dm,500c}$. Similarly, we use the cluster total mass, which is determined by the FoF algorithm. These masses are defined as $M_{\rm bar,FoF}$ for baryons, $M_{\rm dm,FoF}$ for dark matter, and $M_{\rm FoF}\equiv M_{\rm bar,FoF}+M_{\rm dm,FoF}$ for the total mass. The masses measured within $R_{\rm 500c}$ are used to compare scaling relations in simulations with observations, while the FoF masses are employed to examine whether the self-similarity of the scaling relations is satisfied.
@@
For galaxies in clusters, we define two apertures for mass measurement. The first corresponds to the subfind-measured masses of stars, gas (hot, warm, cold, or total), and dark matter. The second corresponds to a spherical aperture with a radius $R_{\rm sph}\equiv 2\times R_{M_{\rm star}/2}$, comparable to the optical radius $R_{\rm sph}\sim r_{\rm op}$, which encloses most of the stellar flux in SED fitting (here, we define $R_{\rm sph}\equiv r_{\rm op}$ to present the results). This latter aperture is beneficial for comparing simulations with observations and to estimate the impact of systematic uncertainties in the scaling relations due to differences in the metrics to measure masses. Finally, we characterize the intracluster medium (ICM) as all the mass within the cluster that do not belong to galaxies according to our definitions. To summarize, Table 2 shows all the mass notations for the study.

### v2

# Differences in baryonic and dark matter scaling relations of galaxy clusters: A comparison between IllustrisTNG simulations and observations

@@
Differences in baryonic and dark matter scaling relations of galaxy clusters: A comparison between IllustrisTNG simulations and observations
@@ 1 Introduction
One of the key discoveries of the last century was that the Universe is expanding. At present, we have a well-constrained standard model, Lambda Cold Dark Matter ($\Lambda$CDM), which states that the Universe is composed of dark energy ($\Omega_{\Lambda}\sim 0.69$) and matter ($\Omega_{\rm m}=\Omega_{\rm b}+\Omega_{\rm dm}\sim 0.31$), where $\Omega_{\rm b}$ represents the baryonic density and $\Omega_{\rm dm}$ denotes the dark matter density. From the density of matter, $15.7\%$ corresponds to baryons, while $\sim$ $84.3\%$ is attributed to dark matter (Planck Collaboration et al., 2016a). This predominance of dark matter over visible matter significantly impacts the gravitational processes driving the formation of structures. Briefly, after the Big Bang, small quantum perturbations in the density field led to matter overdensities. As dark matter only interacts gravitationally, it had time to clump up and grow, forming the first dark matter haloes. These regions started to grow by accreting baryons and smaller haloes (White & Rees, 1978; White & Frenk, 1991). Several authors have thoroughly reviewed this hierarchical structure formation over the last decade, through analytical studies (e.g. Press & Schechter, 1974; White & Rees, 1978; White & Frenk, 1991) and numerical simulations (e.g. White, 1978; Villumsen, 1982; McGee et al., 2009; Pallero et al., 2019).
@@
Within these haloes, galaxy clusters play a key role in helping us to understand the evolution of the Universe, as they represent the largest gravitationally bound structures. To form this kind of structure, matter in large regions of the Universe collapse ($\sim 10$Mpc), resulting in a baryon fraction close to the universal value $f_{\rm bar}^{\rm U}\sim\Omega_{\rm b}/\Omega_{\rm m}\sim 0.157\pm 0.004$ (Planck Collaboration et al., 2016a). Observational results indicate that massive clusters ($\log_{10}{\left(M_{\rm 200c}/\rm M_{\odot}\right)}\gtrsim 14.5$) typically exhibit a baryonic fraction close to this universal value, consistent with expectations (Giodini et al., 2009). Furthermore, Sunyaev-Zel’dovich (SZ) surveys and X-ray measurements confirm similar results (Chiu et al., 2018, C18 hereafter). However, low-mass clusters and groups of galaxies deviate from these expectations (e.g. Ettori, 2003; McGaugh et al., 2010; Lovisari, L. et al., 2015; Chiu et al., 2018).
@@
The last point highlights a discrepancy between early-time observations of baryon fractions in the cosmic microwave background (CMB) (e.g. Planck Collaboration et al., 2016a) and late-time SZ and X-ray observations of galaxy groups and clusters. For example, C18 analysed the baryon fraction, $f_{\rm bar,500c}=f_{\rm star,500c}+f_{\rm hot-gas,500c}$, of 91 self-similar clusters in the redshift range $1.250.20<z<$ 1.25, using data from the SPT-SZ survey and Chandra X-ray observations. Using scaling relations for X-ray hot gas and galaxy stellar mass, they found that clusters with $\log_{10}{\left(M_{\rm 500c}/\rm M_{\odot}\right)}\gtrsim 14.9$ reach a baryon fraction close to the universal value. However, clusters with $\log_{10}{\left(M_{\rm 500c}/\rm M_{\odot}\right)}\lesssim 14.9$ exhibit significantly lower baryon fractions. These observations suggest that not all baryonic matter in low-mass clusters is detected, even though it should reside within their halo boundaries (Shull et al., 2012; Sanderson et al., 2013). This discrepancy presents a key challenge in the $\Lambda$CDM model, commonly termed the missing baryon problem. Additionally, C18 found that baryon fractions vary strongly with total mass but weakly with redshift. These two trends are inconsistent with a hierarchical scenario, as the most massive clusters with baryon fractions near the universal value could not have formed through the accretion of low-mass clusters with lower baryon fractions.
@@
Studies from cosmological simulations and observations led to two main explanations for the missing baryons. The first suggests the existence of warm-hot gas in the intergalactic medium (WHIM) at temperatures of $10^{5}-10^{7}$ K (Cen & Ostriker, 1999). Nicastro et al. (2018) detected diffuse WHIM using two OVII absorption line tracers, although their measurements involve considerable uncertainties, as the signals could originate from sources other than the WHIM itself (Johnson et al., 2019). Furthermore, de Graaff et al. (2019) detected WHIM through the Planck thermal SZ (tSZ) signal (Planck Collaboration et al., 2016b), and WHIM has been characterised and analysed in magnetohydrodynamical simulations (e.g. Martizzi et al., 2019). The second explanation posits that the missing baryons reside in regions outside the halo boundary, such as $R_{500\rm c}$ or $R_{200\rm c}$ (e.g. Haider et al., 2016; Ayromlou et al., 2021; Gouin et al., 2022; Ayromlou et al., 2023). Feedback processes redistribute baryons to the outskirts of the halo, particularly in haloes with $\log_{10}\left(M_{\rm 200c}/M_{\odot}\right)\lesssim 14$ (Ayromlou et al., 2023, A23 hereafter). Using simulations, A23 introduced the closure radius, $R_{\rm c}$, as the radius where the baryon fraction approaches the universal value. For example, for haloes with $\log_{10}\left(M_{\rm 200c}/M_{\odot}\right)\lesssim 14$ , active galactic nucleus (AGN) feedback processes redistribute baryons to $R_{\rm c}\sim 1.5-2.5R_{\rm 200c}$ (Ayromlou et al., 2023). These two explanations are not mutually exclusive and provide complementary insights into galaxy evolution and large-scale structure cosmology (Walker et al., 2019; Nicastro et al., 2022).
@@
Scaling relations are powerful tools for investigating the power laws associated with observables in haloes. The self-similar model (Kaiser, 1986) provides a mathematical framework for describing scaling relations in hydrostatic equilibrium structures. These tools enable the use of large samples of galaxy clusters to constrain cosmological parameters (Haiman et al., 2001; Holder et al., 2001; Carlstrom et al., 2002). Scaling relations have been extensively studied through X-ray observations (e.g. Mohr & Evrard, 1997; Arnaud & Evrard, 1999; Mohr et al., 1999; Reiprich & Böhringer, 2002; O’Hara et al., 2006; Arnaud et al., 2007; Pratt et al., 2009; Sun et al., 2009; Vikhlinin et al., 2009; Mantz et al., 2016; Chiu et al., 2018) and hydrodynamical simulations (e.g. Evrard, 1997; Bryan & Norman, 1998; Nagai et al., 2007; Stanek et al., 2010; Pillepich et al., 2018; Pop et al., 2022a; Hadzhiyska et al., 2023a). The inclusion of AGN feedback processes in simulations significantly affects the X-ray and SZ scaling relations in galaxy groups (Puchwein et al., 2008; Fabjan et al., 2011; Pike et al., 2014; Planelles et al., 2013; Le Brun et al., 2016; Truong et al., 2017; Henden et al., 2018, 2019; Lim et al., 2021; Yang et al., 2022). Pop et al. (2022a, P22 hereafter) introduced a novel method for fitting a scaling relation using a general double power-law (GDPL) model derived from the BPL (Jóhannesson et al., 2006) scheme. The GDPL model demonstrated superior performance in representing self-similarity for both relaxed and unrelaxed haloes at $\log_{10}\left(M_{\rm 500c}/{\rm M_{\odot}}\right)\gtrsim 14$ compared to the simple power law (SPL: Lotka, 1926) in simulations. If the trends and slopes of the SPL scaling relations presented by C18 can be reproduced in simulations, the missing baryons can be identified and further investigated.
@@
In this study, we compared the self-similar scaling relations of galaxy clusters obtained from the observations of C18 with the hydrodynamical simulations. We used the suite of cosmological magnetohydrodynamical simulations from the IllustrisTNG project, selecting a sample of haloes with $M_{\rm 200c}\geq 7\times 10^{13}$ $\rm M_{\odot}$ to match the observational mass conditions and scaling relations of C18. To study their behaviour separately, we introduced cold and warm gas components into these scaling relations. We analysed the percentage distribution of the baryonic components of the haloes: cold gas, WHIM, hot gas, stellar mass from galaxies, and intra-cluster light (ICL). Finally, we fit the SPL, BPL, and GDPL models and examine their respective shapes.
@@
This paper is structured as follows. In Section 2, we describe the methods used in this research, including simulations and observational data. Section 3 outlines the formalism of baryonic self-similar scaling relations, incorporating both the SPL and GDPL approaches. In Section 4, we present the results of the SPL, BPL, and GDPL models for the baryonic fraction scaling relations, comparing these models with observational data, varying the galaxy aperture, and adding the remaining gas components. In Section 5, we discuss the implications of these results, particularly in the context of baryonic component accretion. Finally, in Section 6, we summarise our findings.
@@ 2 Simulations and observational data
2.1 IllustrisTNG: The state-of-the-art magnetohydrodynamical and cosmological simulations
IllustrisTNG (Naiman et al., 2018, Nelson et al., 2018, Springel et al., 2018, Pillepich et al., 2018, Marinacci et al., 2018) is a comprehensive suite of state-of-the-art magnetohydrodynamical and cosmological simulations focused on the formation and evolution of galaxies. It uses the moving-mesh AREPO code (Springel, 2010), which solves the Poisson equations to compute the gravitational potential. In addition, it solves magnetohydrodynamic equations using a Voronoi tessellation method (Pakmor et al., 2011; Pakmor & Springel, 2013). It includes a physical model for galaxy formation, incorporating non-gravitational processes such as gas radiative cooling, star formation, stellar evolution, and stellar feedback (Pillepich et al., 2018). IllustrisTNG also accounts for the physics of supermassive black holes, including their seeding, merging, and feedback processes (Weinberger et al., 2017).
@@ 2.1 IllustrisTNG: The state-of-the-art magnetohydrodynamical and cosmological simulations
We define galaxy clusters as self-gravitating haloes with masses $M_{\rm 200c}\geq 7\times 10^{13}$ $\rm M_{\odot}$ (for any redshift), comparable to the mass of the Fornax cluster (Drinkwater et al., 2001). The haloes and subhaloes were identified in a two-step fashion. First, haloes were selected using the FoF algorithm (FoF: Davis & Djorgovski, 1985; Springel et al., 2001), with a linking length of $b=0.2$, applied only to dark matter particles. Then, subhaloes were selected within FoF haloes using the Subfind algorithm (Springel et al., 2001; Dolag et al., 2009), as those self-gravitating structures residing within haloes. In this work, we define galaxies as those subhaloes containing 1000 stellar particles, corresponding to $M_{\rm star}\geq 5\times 10^{9}$ $\rm M_{\odot}$ for TNG100 and $M_{\rm star}\geq 5\times 10^{10}$ $\rm M_{\odot}$ for TNG300, to ensure sufficient resolution to avoid spurious identifications. The gas within the structures was divided into three components: cold gas ($T<10^{5}$ $\rm K$), warm gas ($10^{5}\leq T<10^{7}$ $\rm K$), and hot gas ($T\geq 10^{7}$ $\rm K$) (Cen & Ostriker, 1999; Martizzi et al., 2019; Gouin et al., 2022, 2023). These temperature thresholds were chosen because the star formation model in IllustrisTNG uses an effective equation of state and a representation of cold and warm gas phases within each cell (Pillepich et al., 2018). Moreover, these thresholds are widely employed in IllustrisTNG studies to distinguish gas phases based on temperature and electron density (e.g. Martizzi et al., 2019; Gouin et al., 2022, 2023), as they effectively capture the gas phase diagram. Clusters were selected within $0.0\leq z\leq 1.5$, with a time window between snapshots of 0.5 Gyr (TNG100) and 1.8 Gyr (TNG300). This resulted in a sample of 218 and 1605 clusters, respectively.
@@
For each cluster, we defined a radius $R_{\Delta}=R_{\rm 500c}$, where the density is $500$ times the critical density of the Universe (see Section 3). The baryon (gas and stars), stellar mass, and the hot, warm and cold gas masses were defined by summing the corresponding particles with $R_{\Delta}$. The total mass, $M_{\rm 500c}$, is the sum of $M_{\rm bar,500c}$ and the dark matter mass, $M_{\rm dm,500c}$. Similarly, we used the cluster total mass, which was determined by the FoF algorithm. These masses were defined as $M_{\rm bar,FoF}$ for baryons, $M_{\rm dm,FoF}$ for dark matter, and $M_{\rm FoF}\equiv M_{\rm bar,FoF}+M_{\rm dm,FoF}$ for the total mass. The masses measured within $R_{\rm 500c}$ were used to compare the scaling relations in the simulations with the observations, while the FoF masses were used to examine whether the self-similarity of the scaling relations was satisfied.
@@
For galaxies in clusters, we defined two apertures for mass measurement. The first corresponded to the Subfind-measured masses of stars, gas (hot, warm, cold, or total), and dark matter. The second corresponded to a spherical aperture with a radius $R_{\rm sph}\equiv 2\times R_{M_{\rm star}/2}$, comparable to the optical radius $R_{\rm sph}\sim r_{\rm op}$, which encloses most of the stellar flux in spectral energy distribution (SED) fitting (hereafter $R_{\rm sph}\equiv r_{\rm op}$). This second aperture was beneficial for comparing simulations with observations and for estimating the impact of systematic uncertainties in the scaling relations due to differences in the metrics to measure masses. Finally, we characterised the intracluster medium (ICM) as all mass within the cluster not belonging to galaxies, as per our definitions. Table 2 summarises all mass notations used in the present study.